OpenStudy (anonymous):

1. Find the gradient of w = x^2y^3z at (1, 2,−1).

OpenStudy (gorv):

grad w= \[\Delta w(x,y,z)=(\frac{ dw }{ dx },\frac{ dw }{ dy },\frac{ dw }{ dz })\]

OpenStudy (gorv):

calculate this and thn place the point (1,2,-1) at place of x , y , z